# Light Rings, Accretion Disks and Shadows of Hayward Boson Stars in asymptotically AdS Spacetime

### v1

@@ 1 Introduction
The general theory of relativity, proposed by Einstein in 1916 Einstein1916 , has been remarkably successful in describing gravitational phenomena, especially in extreme conditions such as strong gravitational fields. Notable examples include the Schwarzschild black hole Schwarzschild1916 , Kerr black hole Kerr1963 and Kerr–Newman black hole NewmanCouchChinnapared1965 . However, the issue of singularities remains unresolved. Singularities are not a reflection of the real physical world but rather a limitation of existing theories when describing the behavior of matter and spacetime under extreme conditions.
@@
To address the singularity problem, regular black holes are proposed as theoretical models to avoid singularity formation. Early attempts to regularize black hole metrics were made by Shirokov (1948) Shirokov1948 , Duan (1954) Duan1954 , Sakharov Sakharov1966 (1966), and Gliner (1966) Gliner1966 . In 1968, Bardeen proposed a widely accepted model of regular black holes Bardeen1968 . Further in 2006, Hayward introduced another model of regular black holes Hayward2006 .
@@
However, obtaining exact solutions for regular black holes by solving the Einstein equations remains highly challenging. It was not until 1999 that Ayón-Beato and García Ayon-BeatoGarcia1998 proposed the first exact regular black hole solution using nonlinear electrodynamics coupled to general relativity. Subsequently, they derived the exact solution for the Bardeen black hole Ayon-BeatoGarcia2000 , interpreting it as a magnetic monopole. Further, the exact solution Hayward black hole is obtained by Fan and Wang in 2016 FanWang2016 .
@@
Building upon the pioneering work of Hawking and Page on black hole thermodynamics in Anti-de Sitter (AdS) spacetime HawkingPage1983 , comprehensive investigations have been carried out in recent years into the thermodynamic properties of the Hayward model in AdS spacetime Fan2017 ; HaldarBiswas2018 ; GuoPuJiang2019 ; TakayanagiTamaoka2020 ; RodrigueSalehThomas2020 ; KumaraRizwanHegde2020 ; GuoHuangHe2021 ; KumaraRizwanHegde2021 ; ZhangZhangZou2021 ; GuoHuangLiang2022 ; BelhajBenaliMoumni2022 ; Iguchi2023 ; RizwanJusufi2023 ; KumaraPunachaHegde2023 ; GuoXieMiao2024 ; GogoiAcharjeePhukon2024 .
@@
Beyond its thermodynamic aspects, the Hayward model has also drawn considerable attention across multiple frontier research areas, including greybody factor and stability analysis WuCaiXie2025 , construction of extended hydrodynamic models NascimentoBezerraToledo2024 , studies on quantum black hole properties CadoniOiSanna2022 , quasinormal modes ToshmatovAbdujabbarovStuchlik2015 ; PedrazaLopezArceo2022 ; RoyKar2022 ; ZhangGongFu2024 ; Malik2024 ; RaoHuang2025 , as well as computations and analyses of black hole shadow radii BelhajBenaliMoumni2022 ; LuoYuCao2023 .
@@
In addition to regular black holes, boson stars are also ideal compact objects that lack singularities and can mimic black holes in many observational features VincentMelianiGrandclement2016 ; CunhaFontHerdeiro2017 ; OlivaresYounsiFromm2020 ; RosaGarciaVincent2022 ; RosaRubiera-Garcia2022 ; RosaMacedoRubiera-Garcia2023 ; RosaPellePerez2024 ; deSaLimaHerdeiro2024 ; GjorgjieskiKunzNedkova2024 ; ZengYangHuang2025 ; LiWuHe2025 ; HeLiYang2025 ; HuangLiuZhang2025 . Studies of boson stars were originally carried out under the assumption of asymptotically flat spacetime RuffiniBonazzola1969 . In recent years, this framework has been extended to asymptotically anti-de Sitter (AdS) spacetimes with a negative cosmological constant AstefaneseiRadu2003 ; HartmannRiedel2013 ; BrihayeHartmannRiedel2015 ; BrihayeConsoleHartmann2022 , and further generalized to other similar models such as Dirac stars ZhangZhaoWang2024 ; ZhangZhaoWang2025 and Proca stars DuarteBrito2016 .
@@
This paper aims to present a analysis of certain fundamental properties of Hayward mini-boson stars in AdS spacetime. Compared to the case with vanishing cosmological constant YueWang2023a , the presence of a negative cosmological constant significantly affects various features of these boson stars, such as their maximum mass, minimum frequency, frozen state, critical horizon radius, and photon sphere radius. In addition, we further investigate the radius, angular momentum, and energy of particles at the innermost stable circular orbit (ISCO) in the accretion disk, and analyze how the negative cosmological constant influences the observable spectral distribution on the accretion disk.
@@
The organization of this paper is as follows. In Sec. 2, the model discussed is presented. In Sec. 3, the numerical scheme is outlined. In Sec. 4, the numerical results and discussion are presented in four topics. In Sec. 5, some conclusions are presented.
2 Introduction to the Model
@@
This section provides a brief introduction to the theoretical model, which combines the Einstein-nonlinear electrodynamics framework proposed in Fan2017 ; FanWang2016 with a complex scalar field. The natural units ($\hbar=c=1$) are adopted, and the action is given by:
@@
where, ^$\hat{\mathcal{F}}=\frac{1}{4}\hat{F}_{\mu\nu}\hat{F}^{\mu\nu}$, $\hat{F}_{\mu\nu}=\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$, $\hat{A}_{\mu}=\{0,0,0,\hat{q}\cos\theta\}$, $\hat{q}=q\sqrt{G}$, $q$ denotes the magnetic monopole charge, and parameter $s$ is free. It is worth noting that the electromagnetic action (2) used in this paper is similar to that in Ayon-BeatoGarcia2000 , but differs from the one adopted in Fan2017 .
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+g_{\mu\nu}\Lambda-2\kappa(T^{(1)}_{\mu\nu}+T^{(2)}_{\mu\nu})$ $=$ $0,$ (5)
$\nabla_{\mu}\left(\frac{\partial\mathcal{L}^{(1)}}{\partial\mathcal{F}}F^{\mu\nu}\right)$ $=$ $0,$ (6)
$\Box\Phi-\dot{U}(\Phi\Phi^{*})\Phi$ $=$ $0$ (7)
@@
$\phi^{\prime\prime}+\left(\frac{2}{r}+\frac{N^{\prime}}{N}+\frac{\sigma^{\prime}}{\sigma}\right)\phi^{\prime}+\left(\frac{\omega^{2}}{N\sigma^{2}}-\frac{\mu^{2}}{2}\right)\frac{\phi}{N}$ $=$ $0,$ (11)
$N^{\prime}+\frac{\kappa\omega^{2}r\phi^{2}}{N\sigma^{2}}+\kappa rN\phi^{\prime 2}+\frac{N-1}{r}+\frac{3\kappa\hat{q}^{6}r}{s(\hat{q}^{3}+r^{3})^{2}}+\left(\kappa\mu^{2}\phi^{2}+\Lambda\right)r$ $=$ $0,$ (12)
$\frac{\sigma^{\prime}}{\sigma}-\kappa r\left(\phi^{\prime 2}+\frac{\omega^{2}\phi^{2}}{N^{2}\sigma^{2}}\right)$ $=$ $0,$ (13)
@@ 3 Numerical scheme
Finally, the boundary value problem defined by equations (11)–(13) and the boundary conditions (14) and (15) is solved using the open-source finite element computing software FEniCSx (version 0.8) BarattaDeanDokken2023 ; ScroggsEtal2022 ; BasixJoss ; AlnaesEtal2014 . We set the computational domain to be discretized with $\geq 10^{4}$ mesh nodes, and the relative numerical error tolerance is set to $<10^{-8}$. The smallest frequency parameter considered in our computations is $\tilde{\omega}=0.0001$. Although solutions for $\tilde{\omega}<0.0001$ could theoretically be obtained by further refining the mesh, numerical convergence becomes increasingly challenging as the stiffness of the system increases.
4 Numerical results
When $\tilde{\Lambda}=0$, numerical solutions to equations (11)–(13) were presented in the literature YueWang2023a . A notable feature of these solutions is that they include a frozen state configuration with a critical event horizon in the limit $\tilde{\omega}\to 0$ (practically down to $\tilde{\omega}=0.0001$). Such solutions endow boson stars with an almost black-hole-like hard boundary, bringing their physical properties closer to those of black holes.
Based on the work presented in YueWang2023a , we aim to further investigate the existence of such frozen states in the case where $\tilde{\Lambda}\neq 0$. Since the existence of frozen state solutions depends on the value of $\tilde{q}$, we begin by determining the minimum value of $\tilde{q}$ for which such solutions can exist. This threshold is referred as the critical value of $q$, and denoted by $\tilde{q}_{c}$.
@@ 4.1 ADM Mass
The increased interval between $\tilde{\Lambda}$ values (compared to Eq. (19)) is selected to avoid overcrowding of the $\tilde{M}$-$\tilde{\omega}$ curves. Figure 3 shows the $\tilde{M}$-$\tilde{\omega}$ curve for four typical values of $\tilde{q}$. As can be seen from Figure 3, as $\tilde{\Lambda}$ increases, the spiral structure disappears, and compared to the case with small $\tilde{q}$, the overall variation range of $\tilde{M}$ is narrower.
It should be noted that for the $\tilde{M}-\tilde{\omega}$ curve without a spiral structure, the value of $\tilde{\omega}$ can continuously decrease (practically down to $\tilde{\omega}=0.0001$). This phenomenon will have a significant impact on the structure of spacetime, which will be discussed in 4.2.
Figure 3: $\tilde{M}$ versus $\tilde{\omega}$ for large $\tilde{q}$ values with $\tilde{s}=0.2$. The minimum of $\tilde{\omega}$ is $0.0001$.
The corresponding dependency of $\tilde{M}_{\mathrm{max}}$ and $\tilde{\omega}_{\mathrm{min}}$ on $\tilde{\Lambda}$ for large $\tilde{q}$ is shown in Fig. 4. Interestingly, $\tilde{M}_{\mathrm{max}}$ now increases monotonically as $\tilde{\Lambda}$ decreases. For $\tilde{\omega}_{\mathrm{min}}$, when its value greater than $0.0001$, it also shows a consistent monotonic increase with decreasing $\tilde{\Lambda}$.
@@
4.2 Frozen States and Critical Horizon
First, the effect of the continuously decreasing $\tilde{\omega}$ on the solutions to the equations of motion is discussed. A set of numerical results for typical $\tilde{q}$ and $\tilde{\Lambda}$ values is displayed in Figure 5, revealing that a significant increase in the absolute values of the $\tilde{\phi}$ and $\sigma$ gradients near $x=0.34$ is induced by reducing $\tilde{\omega}$. This behavior is demonstrated below to have a substantial impact on metrics.
@@ 4.2 Frozen States and Critical Horizon
Figure 6(a) and Figure 6(b) show the variation of $-g_{tt}$ and $g^{rr}$ with $\tilde{\omega}$. It can be observed that as $\tilde{\omega}$ decreases, the minimum values of $-g_{tt}$ and $g^{rr}$ continue to decrease. When $\tilde{\omega}$ approaches $0.0001$, there exists a region centered at $x=0$ where $-g_{tt}$ significantly decreases to the order of $\mathcal{O}(10^{-7})$. Meanwhile, the minimum value of $g^{rr}$ also drops to the same order. If defining the boundary of this region as $x_{c}$, as shown in Figure 6(b), the value of $x_{c}$ can be precisely determined using the minimum value of $g^{rr}$.
@@
As $x\to x_{c}$, $-g_{tt}\to 0$. For a distant observer, time near $x_{c}$ slows down infinitely, causing matter to asymptotically approach—but never cross—the surface at $x_{c}$ within finite time. This behavior aligns with the definition of a "frozen star" ZeldovichNovikov1971 . However, due to the limitations of numerical computation, we cannot obtain a solution where $-g_{tt}|_{x\to x_{c}}\to 0$ in practice. Therefore, we conclude that when $\tilde{\omega}=0.0001$ and $-g_{tt}|_{x\to x_{c}}\approx 10^{-7}$, the boson star is in a "frozen state" YueWang2023a ; HuangSunWang2025 ; ChenWang2024 , and call $x_{c}$ the critical horizon radius. The critical horizon radius $\tilde{r}{c}^{\mathrm{H}}$ in the physical coordinate system is then given by:
@@
Further, Figure 7 illustrates how $\tilde{\Lambda}$ affects both the ADM mass and critical horizon radius $\tilde{r}_{c}^{\mathrm{H}}$ of boson star at frozen state. The ADM mass shows a monotonic increase with decreasing $\tilde{\Lambda}$, while $\tilde{r}_{c}^{\mathrm{H}}$ follows an opposite trend, decreasing monotonically as $\tilde{\Lambda}$ decreasing. If defining $\tilde{r}_{c}^{\mathrm{H}}$ as the effective radius of a boson star, that implies the compactness of the boson star increases as $\tilde{\Lambda}$ decreases.
@@ 4.3 Light Rings
where the dot denotes differentiation with respect to the affine parameter (for photons) or proper time (for massive particles) along the geodesic. Here, $x^{\mu}=(t,\tilde{r},\theta,\varphi)$ represents the coordinate components, with $\epsilon=0$ and $\epsilon=1$ corresponding to photons and massive particles, respectively.
Due to the static spherical symmetry, we assume particles’ orbits lie in the equatorial plane $\theta=\pi/2$, implying $\dot{\theta}=0$. The metric (10) possesses two Killing vectors $\partial_{t}$ and $\partial_{\varphi}$, corresponding to conserved quantities: the energy $E=-g_{tt}\dot{t}$ and angular momentum $L=\tilde{r}^{2}\dot{\varphi}$. Equation (21) then reduces to
@@
The stability of the orbit along the $\theta$-direction of boson stars has been rigorously proved by Cunha, Berti and Herdeiro CunhaBertiHerdeiro2017 .
Figure 8 shows the effective potential curves corresponding to two typical parameter combinations ($\tilde{q}$ and $\tilde{\Lambda}$). From the figure, it can be observed that each effective potential curve has two extreme points, which means there are two light rings (LRs) MacedoPaniCardoso2013 . The positions of the extreme points indicate that the inner light ring is stable while the outer one is unstable.
@@
$\tilde{q}\in\{0.0,0.2,0.4,0.45\},\ \tilde{\Lambda}\in\{0.0,-0.02,-0.04,-0.06,-0.08,-0.1\},$
@@
$\tilde{q}\in\{0.5,0.52,0.54,0.56,0.58,0.60\},$
$\tilde{\Lambda}\in\{0.0,-0.2,-0.4,-0.5,-0.6,-0.8,-1.0,-1.5,-2.0,-2.5,-3.0\}.$
@@
Further, the $\tilde{\omega}$ dependence of these LR radii is presented in Figure 11 for some large values of $\tilde{q}$ and a broader range of $\tilde{\Lambda}$ values, where the inner and outer radii are denoted by $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$, respectively. From the figure, it is found that: (1) as $\tilde{\Lambda}$ is decreased, the relationship between $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$ is changed from linear to nonlinear; and (2) as $\tilde{\omega}$ is increased, the radii $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$ are brought increasingly close.
@@
To intuitively demonstrate the quantitative relationship between the LR radii and the critical horizon radius of frozen boson stars, as well as their dependence on $\tilde{\Lambda}$, their numerical values are presented in Figure 12. It is shown that both the inner LR radius and critical horizon radius are increased monotonically as $\tilde{\Lambda}$ is decreased, while the outer LR radius is observed to exhibit an opposite trend, being decreased monotonically with $\tilde{\Lambda}$. Furthermore, the inner light ring is always located within the critical horizon ($\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}<x_{c}$), while the outer light ring is consistently positioned outside this boundary ($\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}>x_{c}$).
@@
To determine photon trajectories orbiting a photon sphere, we numerically solve Equation (22). Introducing the substitution $u=1/\tilde{r}$ Luminet1979 and using angular momentum conservation ($L=\tilde{r}^{2}\dot{\varphi}$), Equation (22) may be transformed to:
@@
Figure 13 illustrates a typical halo image obtained through numerical integration. The gray disk in the figure corresponds to the region within the critical horizon radius, while the blue ring depicts the trajectory of photons completing two full orbits. The red marker indicates the emission point, where photons are emitted vertically upward.
The images reveal that as the parameter $\tilde{\Lambda}$ decreases, both the critical horizon radius and the inner light ring shrink, whereas the outer light ring exhibits a clear expansion. These phenomena are consistent with the results presented in Figure 12.
@@
4.4 Accretion Disk and Shadows
Here we assume that (1) the accretion disk is geometrically thin and optically thick NovikovThorne1973 ; PageThorne1974 , and (2) the observer’s reference frame is at rest relative to the boson star, satisfying the zero angular momentum observer (ZAMO) condition.

### v2

@@ I Introduction
The general theory of relativity, proposed by Einstein in 1916 [1], has been remarkably successful in describing gravitational phenomena, especially in extreme conditions such as strong gravitational fields. Notable examples include the Schwarzschild black hole [2], Kerr black hole [3] and Kerr–Newman black hole [4]. However, the issue of singularities remains unresolved. Singularities are not a reflection of the real physical world but rather a limitation of existing theories when describing the behavior of matter and spacetime under extreme conditions.
@@
To address the singularity problem, regular black holes are proposed as theoretical models to avoid singularity formation. Early attempts to regularize black hole metrics were made by Shirokov (1948) [5], Duan (1954) [6], Sakharov [7] (1966), and Gliner (1966) [8]. In 1968, Bardeen proposed a widely accepted model of regular black holes [9]. Further in 2006, Hayward introduced another model of regular black holes [10].
@@
However, obtaining exact solutions for regular black holes by solving the Einstein equations remains highly challenging. It was not until 1999 that Ayón-Beato and García [11] proposed the first exact regular black hole solution using nonlinear electrodynamics coupled to general relativity. Subsequently, they derived the exact solution for the Bardeen black hole [12], interpreting it as a magnetic monopole. Further, the exact solution Hayward black hole is obtained by Fan and Wang in 2016 [13].
@@
Building upon the pioneering work of Hawking and Page on black hole thermodynamics in Anti-de Sitter (AdS) spacetime [14], comprehensive investigations have been carried out in recent years into the thermodynamic properties of the Hayward model in AdS spacetime [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@
Beyond its thermodynamic aspects, the Hayward model has also drawn considerable attention across multiple frontier research areas, including greybody factor and stability analysis [31], construction of extended hydrodynamic models [32], studies on quantum black hole properties [33], quasinormal modes [34, 35, 36, 37, 38, 39, 40], black hole shadow radii [25, 41, 42], as well as primordial regular black holes as the dark matter [43]
@@
In addition to regular black holes, boson stars are also ideal compact objects that lack singularities and can mimic black holes in many observational features [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Studies of boson stars were originally carried out under the assumption of asymptotically flat spacetime [57]. In recent years, this framework has been extended to asymptotically anti-de Sitter (AdS) spacetimes with a negative cosmological constant [58, 59, 60, 61], and further generalized to other similar models such as Dirac stars [62, 63] and Proca stars [64].
@@
This paper aims to present a analysis of certain fundamental properties of Hayward mini-boson stars in AdS spacetime. Compared to the case with vanishing cosmological constant [65], the presence of a negative cosmological constant significantly affects various features of these boson stars, such as their maximum mass, minimum frequency, frozen state, critical horizon radius, and photon sphere radius. In addition, we further investigate the radius, angular momentum, and energy of particles at the innermost stable circular orbit (ISCO) in the accretion disk, and analyze how the negative cosmological constant influences the observable spectral distribution on the accretion disk.
@@
The organization of this paper is as follows. In Sec. II, the model discussed is presented. In Sec. III, the numerical scheme is outlined. In Sec. IV, the numerical results and discussion are presented in four topics. In Sec. V, some conclusions are presented.
II Introduction to the Model
@@
This section provides a brief introduction to the theoretical model, which combines the Einstein-nonlinear electrodynamics framework proposed in [15, 13] with a complex scalar field. The natural units ($\hbar=c=1$) are adopted, and the action is given by:
@@
where, ^$\hat{\mathcal{F}}=\frac{1}{4}\hat{F}_{\mu\nu}\hat{F}^{\mu\nu}$, $\hat{F}_{\mu\nu}=\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$, $\hat{A}_{\mu}=\{0,0,0,\hat{q}\cos\theta\}$, $\hat{q}=q\sqrt{G}$, $q$ denotes the magnetic monopole charge, and parameter $s$ is free. It is worth noting that the electromagnetic action (2) used in this paper is similar to that in [12], but differs from the one adopted in [15].
@@
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R+g_{\mu\nu}\Lambda-2\kappa(T^{(1)}_{\mu\nu}+T^{(2)}_{\mu\nu})$ $=$ 0 0,0 , (5)
$\nabla_{\mu}\left(\frac{\partial\mathcal{L}^{(1)}}{\partial\mathcal{F}}F^{\mu\nu}\right)$ $=$ 0 0,0 , (6)
$\Box\Phi-\dot{U}(\Phi\Phi^{*})\Phi$ $=$ 0 0.0 . (7)
@@
$\phi^{\prime\prime}+\left(\frac{2}{r}+\frac{N^{\prime}}{N}+\frac{\sigma^{\prime}}{\sigma}\right)\phi^{\prime}+\left(\frac{\omega^{2}}{N\sigma^{2}}-\frac{\mu^{2}}{2}\right)\frac{\phi}{N}$ $=$ 0 0,0 , (11)
$N^{\prime}+\frac{\kappa\omega^{2}r\phi^{2}}{N\sigma^{2}}+\kappa rN\phi^{\prime 2}+\frac{N-1}{r}+\frac{3\kappa\hat{q}^{6}r}{s(\hat{q}^{3}+r^{3})^{2}}+\left(\kappa\mu^{2}\phi^{2}+\Lambda\right)r$ $=$ 0 0,0 , (12)
$\frac{\sigma^{\prime}}{\sigma}-\kappa r\left(\phi^{\prime 2}+\frac{\omega^{2}\phi^{2}}{N^{2}\sigma^{2}}\right)$ $=$ 0 0,0 , (13)
@@ III Numerical scheme
Finally, the boundary value problem defined by equations (11)–(13) and the boundary conditions (14) and (15) is solved using the open-source finite element computing software FEniCSx (version 0.8) [66, 67, 68, 69]. We set the computational domain to be discretized with $\geq 10^{4}$ mesh nodes, and the relative numerical error tolerance is set to $<10^{-8}$. The smallest frequency parameter considered in our computations is $\tilde{\omega}=0.0001$. Although solutions for $\tilde{\omega}<0.0001$ could theoretically be obtained by further refining the mesh, numerical convergence becomes increasingly challenging as the stiffness of the system increases.
IV Numerical results
When $\tilde{\Lambda}=0$, numerical solutions to equations (11)–(13) were presented in the literature [65]. A notable feature of these solutions is that they include a frozen state configuration with a critical event horizon in the limit $\tilde{\omega}\to 0$ (practically down to $\tilde{\omega}=0.0001$). Such solutions endow boson stars with an almost black-hole-like hard boundary, bringing their physical properties closer to those of black holes.
Based on the work presented in [65], we aim to further investigate the existence of such frozen states in the case where $\tilde{\Lambda}\neq 0$. Since the existence of frozen state solutions depends on the value of $\tilde{q}$, we begin by determining the minimum value of $\tilde{q}$ for which such solutions can exist. This threshold is referred as the critical value of $q$, and denoted by $\tilde{q}_{c}$.
@@ IV.1 ADM Mass
The increased interval between $\tilde{\Lambda}$ values (compared to Eq. (19)) is selected to avoid overcrowding of the $\tilde{M}$-$\tilde{\omega}$ curves. Fig. 3 shows the $\tilde{M}$-$\tilde{\omega}$ curve for four typical values of $\tilde{q}$. As can be seen from Fig. 3, as $\tilde{\Lambda}$ increases, the spiral structure disappears, and compared to the case with small $\tilde{q}$, the overall variation range of $\tilde{M}$ is narrower.
It should be noted that for the $\tilde{M}-\tilde{\omega}$ curve without a spiral structure, the value of $\tilde{\omega}$ can continuously decrease (practically down to $\tilde{\omega}=0.0001$). This phenomenon will have a significant impact on the structure of spacetime, which will be discussed in subsection  IV.2.
Figure 3: $\tilde{M}$ versus $\tilde{\omega}$ for large $\tilde{q}$ values with $\tilde{s}=0.2$. The minimum of $\tilde{\omega}$ is 0.00010.00010.0001.
The corresponding dependency of $\tilde{M}_{\mathrm{max}}$ and $\tilde{\omega}_{\mathrm{min}}$ on $\tilde{\Lambda}$ for large $\tilde{q}$ is shown in Fig. 4. Interestingly, $\tilde{M}_{\mathrm{max}}$ now increases monotonically as $\tilde{\Lambda}$ decreases. For $\tilde{\omega}_{\mathrm{min}}$, when its value greater than 0.00010.00010.0001, it also shows a consistent monotonic increase with decreasing $\tilde{\Lambda}$.
@@
IV.2 Frozen States and Critical Horizon
First, the effect of the continuously decreasing $\tilde{\omega}$ on the solutions to the equations of motion is discussed. A set of numerical results for typical $\tilde{q}$ and $\tilde{\Lambda}$ values is displayed in Fig. 5, revealing that a significant increase in the absolute values of the $\tilde{\phi}$ and $\sigma$ gradients near $x=0.34$ is induced by reducing $\tilde{\omega}$. This behavior is demonstrated below to have a substantial impact on metrics.
@@ IV.2 Frozen States and Critical Horizon
Fig. 6(a) and Fig. 6(b) show the variation of $-g_{tt}$ and $g^{rr}$ with $\tilde{\omega}$. It can be observed that as $\tilde{\omega}$ decreases, the minimum values of $-g_{tt}$ and $g^{rr}$ continue to decrease. When $\tilde{\omega}$ approaches 0.00010.00010.0001, there exists a region centered at $x=0$ where $-g_{tt}$ significantly decreases to the order of $\mathcal{O}(10^{-7})$. Meanwhile, the minimum value of $g^{rr}$ also drops to the same order. If defining the boundary of this region as $x_{c}$, as shown in Fig. 6(b), the value of $x_{c}$ can be precisely determined using the minimum value of $g^{rr}$.
@@
As $x\to x_{c}$, $-g_{tt}\to 0$. For a distant observer, time near $x_{c}$ slows down infinitely, causing matter to asymptotically approach—but never cross—the surface at $x_{c}$ within finite time. This behavior aligns with the definition of a ”frozen star” [70]. However, due to the limitations of numerical computation, we cannot obtain a solution where $-g_{tt}|_{x\to x_{c}}\to 0$ in practice. Therefore, we conclude that when $\tilde{\omega}=0.0001$ and $-g_{tt}|_{x\to x_{c}}\approx 10^{-7}$, the boson star is in a ”frozen state” [65, 71, 72], and call $x_{c}$ the critical horizon radius. The critical horizon radius $\tilde{r}{c}^{\mathrm{H}}$ in the physical coordinate system is then given by:
@@
Further, Fig. 7 illustrates how $\tilde{\Lambda}$ affects both the ADM mass and critical horizon radius $\tilde{r}_{c}^{\mathrm{H}}$ of boson star at frozen state. The ADM mass shows a monotonic increase with decreasing $\tilde{\Lambda}$, while $\tilde{r}_{c}^{\mathrm{H}}$ follows an opposite trend, decreasing monotonically as $\tilde{\Lambda}$ decreasing. If defining $\tilde{r}_{c}^{\mathrm{H}}$ as the effective radius of a boson star, that implies the compactness of the boson star increases as $\tilde{\Lambda}$ decreases.
@@ IV.3 Light Rings
where the dot denotes differentiation with respect to the affine parameter (for photons) or proper time (for massive particles) along the geodesic. Here, $x^{\mu}=(\tilde{t},\tilde{r},\theta,\varphi)$ represent the coordinate components, with $\tilde{t}=t/\mu$. The parameter $\epsilon=0$ corresponds to photons, and $\epsilon=1$ corresponds to massive particles.
Due to the static spherical symmetry, we assume particles’ orbits lie in the equatorial plane $\theta=\pi/2$, implying $\dot{\theta}=0$. The metric (10) possesses two Killing vectors $\partial_{\tilde{t}}$ and $\partial_{\varphi}$, corresponding to conserved quantities: the energy $E=-g_{tt}\dot{\tilde{t}}$ and angular momentum $L=\tilde{r}^{2}\dot{\varphi}$. Equation (21) then reduces to
@@
The stability of the orbit along the $\theta$-direction of boson stars has been rigorously proved by Cunha, Berti and Herdeiro [73].
Fig. 8 shows the effective potential curves corresponding to two typical parameter combinations ($\tilde{q}$ and $\tilde{\Lambda}$). From the figure, it can be observed that each effective potential curve has two extreme points, which means there are two light rings (LRs) [74]. The positions of the extreme points indicate that the inner light ring is stable while the outer one is unstable.
@@
$\tilde{q}\in\{0.0,0.2,0.4,0.45\},\ \tilde{\Lambda}\in\{0.0,-0.02,-0.04,$
0.06,0.08,0.1},-0.06,-0.08,-0.1\},- 0.06 , - 0.08 , - 0.1 } ,
@@
$\tilde{q}\in\{0.5,0.52,0.54,0.56,0.58,0.60\},\ \tilde{\Lambda}\in\{0.0,-0.2,$
0.4,0.5,0.6,0.8,1.0,1.5,2.0,2.5,3.0}.-0.4,-0.5,-0.6,-0.8,-1.0,-1.5,-2.0,-2.5,-3.0\}.- 0.4 , - 0.5 , - 0.6 , - 0.8 , - 1.0 , - 1.5 , - 2.0 , - 2.5 , - 3.0 } .
@@
Further, the $\tilde{\omega}$ dependence of these LR radii is presented in Fig. 11 for some large values of $\tilde{q}$ and a broader range of $\tilde{\Lambda}$ values, where the inner and outer radii are denoted by $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$, respectively. From the figure, it is found that: (1) as $\tilde{\Lambda}$ is decreased, the relationship between $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$ is changed from linear to nonlinear; and (2) as $\tilde{\omega}$ is increased, the radii $\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}$ and $\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}$ are brought increasingly close.
@@
To intuitively demonstrate the quantitative relationship between the LR radii and the critical horizon radius of frozen boson stars, as well as their dependence on $\tilde{\Lambda}$, their numerical values are presented in Fig. 12. It is shown that both the inner LR radius and critical horizon radius are increased monotonically as $\tilde{\Lambda}$ is decreased, while the outer LR radius is observed to exhibit an opposite trend, being decreased monotonically with $\tilde{\Lambda}$. Furthermore, the inner light ring is always located within the critical horizon ($\tilde{r}_{\mathrm{inner}}^{\mathrm{LR}}<x_{c}$), while the outer light ring is consistently positioned outside this boundary ($\tilde{r}_{\mathrm{outer}}^{\mathrm{LR}}>x_{c}$).
@@
To determine photon trajectories orbiting a photon sphere, we numerically solve Equation (22). Introducing the substitution $u=1/\tilde{r}$ [75] and using angular momentum conservation, Equation (22) may be transformed to:
@@
Fig. 13 illustrates a typical halo image obtained through numerical integration. The gray disk in the figure corresponds to the region within the critical horizon radius, while the blue ring depicts the trajectory of photons completing two full orbits. The red marker indicates the emission point, where photons are emitted vertically upward.
The images reveal that as the parameter $\tilde{\Lambda}$ decreases, both the critical horizon radius and the inner light ring shrink, whereas the outer light ring exhibits a clear expansion. These phenomena are consistent with the results presented in Fig. 12.
@@
IV.4 Accretion Disk and Shadows
Here we assume that (1) the accretion disk is geometrically thin and optically thick [76, 77], and (2) the observer’s reference frame is at rest relative to the boson star, satisfying the zero angular momentum observer (ZAMO) condition.